# AweDist: Attention-aware Embedding Distillation for New Input Token Embeddings

### v1

@@ 5 Results
Mistral-7B Llama-3-8B Llama-3-8B-i Llama-3.1-8B Llama-3.1-8B-i Llama-3.2-3B Llama-3.2-3B-i OLMo-2-7B-i Avg. Target 64.5 69.8 70.6 69.2 72.3 60.1 64.4 61.3 66.5 Random (Hewitt, 2021) 57.0±0.6 58.8±0.5 60.6±0.3 59.3±0.5 62.6±0.6 51.6±0.4 55.8±0.2 53.9±0.6 57.5 NTP 55.7±0.5 63.8±0.1 63.8±0.2 62.9±0.4 66.8±0.4 51.5±0.6 55.5±0.5 58.0±0.3 59.8 Mean (Gee et al., 2022) 58.3 63.8 63.4 63.7 66.9 54.2 58.4 58.0 60.8 NTP++ (Lampinen and McClelland, 2018) 61.2±0.4 65.6±0.3 65.3±0.6 66.0±0.5 68.9±0.2 56.6±0.6 61.3±0.5 58.7±0.6 63.0 ZeTT (Minixhofer et al., 2024) 62.7 66.1 66.3 – – – – – – AweDist-NTP++ 63.0±0.5 67.2±0.3 66.7±1.6 67.2±0.3 70.6±0.4 57.6±0.3 62.2±0.3 59.8±0.5 64.3 AweDist 62.8±0.5 67.3±0.2 67.6±0.3 67.3±0.5 71.0±0.2 56.2±1.9 63.1±0.2 61.2±0.3 64.6 AweDist-NTP++ 62.8±0.5 67.6±0.4 67.8±0.5 67.4±0.4 70.9±0.3 57.9±0.1 62.5±0.6 61.2±0.2 64.7
Method Mistral-7B Llama-3-8B Llama-3-8B-i Llama-3.1-8B Llama-3.1-8B-i Llama-3.2-3B Llama-3.2-3B-i OLMo-7B-i Avg. Sim Corr Sim Corr Sim Corr Sim Corr Sim Corr Sim Corr Sim Corr Sim Corr Sim Corr Target 99.8 96.5 100.0 94.3 100.0 98.4 100.0 94.7 100.0 98.4 100.0 93.2 100.0 93.8 99.4 96.9 99.9 95.8 Random (Hewitt, 2021) 0.0 0.4 0.0 0.0 0.0 0.2 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.1 Mean (Gee et al., 2022) 2.5 4.3 16.8 16.2 17.6 20.7 25.6 25.6 23.2 28.3 15.2 15.2 12.1 14.6 20.1 24.2 16.6 18.6 NTP 26.8 30.3 33.2 39.5 34.4 43.4 28.5 35.0 41.4 50.0 19.7 20.7 18.6 21.1 22.5 26.6 28.1 33.3 NTP++ (Lampinen and McClelland, 2018) 49.8 58.2 45.3 49.6 48.4 58.0 58.6 65.6 60.2 69.9 51.6 58.4 50.0 58.8 52.1 57.0 52.0 59.4 ZeTT (Minixhofer et al., 2024) 63.5 68.6 69.5 73.6 70.7 80.3 – – – – – – – – – – – – AweDist 79.7 85.2 72.7 76.2 79.7 91.0 75.8 80.1 81.6 89.8 0.0 0.0 75.0 83.0 83.8 89.5 68.5 74.4 AweDist-NTP++ 77.3 82.6 68.4 73.8 75.8 86.3 72.3 79.9 77.3 86.1 62.3 69.1 65.2 75.8 65.4 74.6 70.5 78.5 AweDist-NTP++ 76.8 82.0 71.7 75.2 80.9 91.6 77.5 81.2 81.1 89.5 72.7 79.5 71.9 80.5 81.2 86.9 76.7 83.3
@@ 5.1 Main Experiments
We provide our main results on benchmarks in Table 1 (reporting a macro-average over benchmark tasks) and our definition generation results in Table 2 (reporting similarity to the original target model’s generation and correctness as judged by a larger LLM). On average, our method outperforms all other baseline embedding initialization methods. In particular, AweDist shows superior results on benchmarks and in its ability to generate correct definitions, which is indicative of higher-quality representations. Additionally, AweDist also exhibits greater similarity to the original tokenization behavior (“Target”) than all other methods, which substantiates our claim that a distillation-based objective is able to better approximate complex multi-subtoken interactions into a single token embedding.
@@
Note that even with its very competitive performance, AweDist does not fully attain the level of the original tokenization, which benefits from the massive pretraining of the original model. This is expected in our zero-shot tokenizer transfer setting without any further training (see Minixhofer et al., 2024). We seek to directly investigate the embedding initialization quality and the advantages of our distillation-based objective compared to next-token prediction rather than the effects of further training. However, results can naturally be improved via further training; we argue that AweDist will provide the best starting point for such further processing. In the following, we analyze various aspects of the results in further detail.
@@
Note that in Table 2, only for Llama-3.2-3B, AweDist has performance on par with random initialization. Llama-3.2-3B(-i) are the only models in our lineup with tied embedding weights. Our objective does not explicitly enforce a bound on the norm of the new embedding, which in this case led to the failure mode of always generating a specific new embedding with very large norm. Note that this does not always happen for checkpoints with tied weights, as evidenced by the results for Llama-3.2-3B-i. However, we additionally propose a simple modification to alleviate the occurrence of this issue by combining AweDist with the next-token prediction objective, which (implicitly) acts as a regularizer on the output embedding norm. In support of our argument that a subtoken attention distillation objective is superior to next-token prediction for learning new token embeddings, the combination via a sum of the two objectives (AweDist-NTP++) generally yields worse results than AweDist alone. However, we can add a dynamic downweighting factor (see Section 4.2) to the next-token prediction objective (AweDist-NTP++), which mostly alleviates the negative interference while keeping the regularizing effect.
@@
Vanilla NTP-tuning of the embeddings yields disappointing results, even underperforming the subtoken mean initialization. We investigate this and observe a degradation of the original token embeddings. Note that for AweDist, we only optimize new token embeddings. Therefore, we also compare against an improved NTP baseline (NTP++) where we similarly optimize only the new token embeddings. NTP++ outperforms NTP, supporting our analysis. However, even with this improvement, AweDist in turn still further surpasses NTP++.
We note that masking gradient updates to the original embeddings during further training is not commonly done when adding new token embeddings, even during initial “embedding tuning” phases where all weights but the embedding layers are frozen. In these phases, the goal is to improve the initialized embeddings for new tokens or “adapt” the new embedding matrix to the existing Transformer layers (de Vries and Nissim, 2021). When normally sampling from a large-scale training corpus, the degradation of original embeddings will be less dramatic, as the number of new tokens per batch will be much lower. However, this is much less computationally efficient and – at least initially – still potentially harmful. We leave a further investigation of this as a practically useful direction for future work.
@@
ZeTT, which uses a pretrained hypernetwork, is the strongest baseline we compare against. We only compare against ZeTT when a pretrained hyper-network by Minixhofer et al. (2024) is available555Minixhofer et al. (2024) do not provide a checkpoint trained specifically for the -instruct variant of Llama-3-8B but show that a transfer from the base version is possible.. We first note that ZeTT yields quite impressive results, outperforming even our optimized NTP++ baseline. Evidently, the pretraining of ZeTT’s hypernetwork has internalized a better embedding prediction than the iterative gradient descent optimization on selected samples we employ for NTP++. Nevertheless, our proposed method AweDist outperforms even ZeTT – without needing any hyper-network pretraining. It is important to note that at inference time, hyper-network based methods such as ZeTT are actually faster than AweDist (and NTP-based baselines), since only a single forward pass through the hyper-network per token is required. However, they require expensive pretraining for each new model and – in some cases – target domain. Also, we can actually use ZeTT-generated embeddings as a starting point and further tune them using our method for even better results. We analyze this possibility in Section 5.2.
@@ 5.2 In-Depth Analysis
Mistral-7B Llama-3-8B Llama-3-8B-i Llama-3.1-8B Llama-3.1-8B-i Llama-3.2-3B Llama-3.2-3B-i OLMo-2-7B-i Avg. Target 93.4 88.1 99.2 91.1 97.8 65.2 90.9 94.0 90.0 Random (Hewitt, 2021) 0.0 0.0 0.0 0.0 0.0 0.0 0.2 0.0 0.0 Mean (Gee et al., 2022) 2.2 4.2 9.1 8.0 11.3 4.2 7.4 7.4 6.7 NTP++ (Lampinen and McClelland, 2018) 34.0 54.3 70.6 49.7 71.6 29.0 53.9 70.0 54.1 ZeTT (Minixhofer et al., 2024) 15.3 25.4 28.4 – – – – – – AweDist 46.3 68.0 86.3 61.8 81.3 31.6 57.5 84.3 64.6 AweDist-NTP++ 48.3 68.0 85.3 60.2 79.7 38.2 63.4 80.9 65.5 AweDist-NTP++ 46.9 71.0 85.7 62.0 80.1 37.2 66.4 83.7 66.6
Mistral-7B Llama-3-8B Llama-3-8B-i Llama-3.1-8B Llama-3.1-8B-i Llama-3.2-3B Llama-3.2-3B-i OLMo-2-7B-i Avg. AweDist (generated data) 62.4 67.1 67.6 67.0 70.1 57.1 62.3 60.9 64.3 AweDist 62.8±0.5 67.3±0.2 67.6±0.3 67.3±0.5 71.0±0.2 56.2±1.9 63.1±0.2 61.2±0.3 64.6
Mistral-7B Llama-3-8B Llama-3-8B-i Llama-3.1-8B Llama-3.1-8B-i Llama-3.2-3B Llama-3.2-3B-i OLMo-2-7B-i Avg. AweDist-Logits 63.0±0.4 67.2±0.3 67.5±0.8 67.3±0.2 70.7±0.5 57.4±0.3 62.8±0.7 60.2±0.2 64.5 AweDist-NTP++ 63.0±0.5 67.2±0.3 66.7±1.6 67.2±0.3 70.6±0.4 57.6±0.3 62.2±0.3 59.8±0.5 64.3 AweDist-NTP++ 62.8±0.5 67.6±0.4 67.8±0.5 67.4±0.4 70.9±0.3 57.9±0.1 62.5±0.6 61.2±0.2 64.7 AweDist 62.8±0.5 67.3±0.2 67.6±0.3 67.3±0.5 71.0±0.2 56.2±1.9 63.1±0.2 61.2±0.3 64.6 AweDist-KL-NTP++ 62.3±0.4 66.7±0.2 66.4±0.3 66.7±0.7 70.1±0.4 57.3±0.2 62.1±0.6 59.7±0.3 63.9 AweDist-KL-NTP++ 62.8±0.5 67.1±0.4 66.3±1.2 67.1±0.2 70.5±0.4 57.5±0.5 62.5±0.2 60.5±0.5 64.3 AweDist-KL 63.0±0.2 67.3±0.2 67.2±0.2 67.1±0.2 70.8±0.5 57.5±0.4 62.7±0.3 60.5±0.6 64.5
@@
In Table 3, we report LLM-as-a-Judge results on the correctness of generated definitions. This tests the limits of our method, as the new tokens are now more complicated, such as "Software Development" or even common phrases such as "spill the beans". In this setting, even the original target model sometimes is not able to generate correct definitions. In turn, the gap of AweDist to the target model is also larger in this experiment. However, AweDist still outperforms all baselines by a large margin. ZeTT (Minixhofer et al., 2024) in particular struggles in the multi-word token setup, likely because it is out-of-distribution from the hyper-network pretraining.
@@
Choice of initial embedding for AweDist.
We find that the subtoken mean initialization as starting point yields better results than random initialization (see “Random + AweDist” in Table 6). If a pretrained embedding prediction hyper-network is available, we can instead use these predicted embeddings as the starting point. We run this experiment using ZeTT; this combination further improves performance compared to using the subtoken mean (see Table 6). We also compare refining the hypernetwork embeddings with NTP++. This greatly improves over NTP++ but yields only minimal gains over using the hypernetwork embeddings themselves without further NTP++.
@@
Our main proposed method AweDist computes the distillation objective based on a mean-squared error (MSE) of last layer hidden states. A natural question is whether we can also use the “traditional” knowledge distillation objective (Hinton et al., 2015) based on the Kullback-Leibler divergence (KL-Div) (Kullback and Leibler, 1951). We analyze this and report the results in Table 5. We additionally include computing the MSE between logits instead of hidden states. For the KL-Div based results, we include different variations of combining the objective with a NTP-based objective. In general, the choice of distillation objective does not matter much. This is hardly surprising, as the logits are merely a linear projection (potentially with an additional normalization layer) of the hidden state; the log-probabilities used for KL-Div are then in turn merely an additional log-normalization via log-softmax. Combining the KL-based results with NTP-based objectives follows similar trends as their combinations with hidden state MSE-based objectives. We opt for the MSE on hidden states for two reasons: (1) In terms of implementation convenience, the objectives with a projection to logits over the vocabulary require a masking step for the “student” model, as the original “teacher” model does not have the new tokens in its vocabulary. (2) Choosing hidden states allows us to explore other layers than just the last one with potentially large computational speedups – we explore this next.
Mistral-7B Llama-3-8B Llama-3-8B-i Avg. Target 64.5 69.8 70.6 68.3 Random (Hewitt, 2021) 56.7 58.7 60.6 58.6 Mean (Gee et al., 2022) 58.3 63.8 63.4 61.8 ZeTT (Minixhofer et al., 2024) 62.7 66.1 66.3 65.0 (Mean +) NTP++ 61.5 65.9 65.4 64.3 ZeTT + NTP++ 62.8 66.1 66.8 65.3 Random + AweDist 62.5 65.3 65.9 64.6 (Mean +) AweDist 63.5 67.3 68.0 66.3 ZeTT + AweDist 64.2 67.9 68.4 66.8 (Mean +) AweDist-NTP++ 63.2 67.2 68.4 66.3 ZeTT + AweDist-NTP++ 64.1 67.9 68.7 66.9
@@
In our main experiments, we apply the AweDist objective on the hidden states after the last layer. However, the last hidden state might not be the most optimal, as it might be overspecialized for next-token prediction (Rogers et al., 2020). In Figure 2, we plot the average performance on our biomedical benchmark datasets using Llama-3-8B-Instruct while varying the target layer for extracting hidden states for our AweDist objective. Indeed we do see a slight downward trend as we approach the last layer – choosing a different layer than the last one can further boost our already strong results. We repeat the analysis with Llama-3.1-8B-Instruct in Appendix C with matching results.
Interestingly, good results can already be achieved with very early layers, e.g., after layer four. This suggests that much of the subtoken-specific contextual information is already added to the residual stream in early layers, which echoes findings of previous work (Vulić et al., 2020; Elhage et al., 2022; Lad et al., 2024; Nakash et al., 2025). This opens up an avenue to further speed up our method by terminating the forward pass after our target layer (and also saving that compute on the backward pass), which can be much faster if we select early target layers. We leave further exploration of this as an exciting direction for future work. For our main experiments, we choose to keep the last layer because this choice does not necessitate a model-specific sweep over target layers. Also, the last layer is a principled choice, as it guarantees that no subtoken interactions that are only modeled in later layers are excluded from the objective.

### v2

@@ 5.1 Main Experiments
We provide our main results on benchmarks in Table 1 (biomedical domain adaptation) and Table 3 (French language adaptation). In Table 2, we report our experiments on definition generation for newly added biomedical tokens (reporting similarity to the original target model’s generation and correctness as judged by a larger LLM). On average, our method outperforms all other baseline embedding initialization methods. In particular, Token Distillation shows superior results on benchmarks and in its ability to generate correct definitions, which is indicative of higher-quality representations. Additionally, Token Distillation also exhibits greater similarity to the original tokenization behavior than all other methods, which substantiates our claim that a distillation-based objective is able to better approximate complex multi-subtoken interactions into a single token embedding.
@@
Note that even with its very competitive performance, Token Distillation does not fully attain the level of the original tokenization, which benefits from the massive pretraining of the original model. This is expected in our zero-shot tokenizer transfer setting without any further training (see Minixhofer et al., 2024), although remarkably, Token Distillation is able to outperform the original tokenization when adding French tokens to Llama3-8B-i without any training of the Transformer layers. We seek to directly investigate the embedding initialization quality and the advantages of our distillation-based objective compared to next-token prediction rather than the effects of further training. However, results can naturally be improved via further training; we argue that Token Distillation will provide the best starting point for such further processing. Additionally, we show in Appendix C that an additional epoch of further Token Distillation is able to improve results, whereas an additional epoch with the next-token prediction objective plateaus. In the following, we analyze various aspects of the results in further detail.
@@
Note that in Table 2, only for Llama3.2-3B, Token Distillation has performance on par with random initialization. Llama3.2-3B(-i) are the only models in our lineup with tied embedding weights. Our objective does not explicitly enforce a bound on the norm of the new embedding, which in this case led to the failure mode of always generating a specific new embedding with very large norm. Note that this does not always happen for checkpoints with tied weights, as evidenced by the results for Llama3.2-3B-i. However, we additionally propose a simple modification to alleviate the occurrence of this issue by combining Token Distillation with the next-token prediction objective, which (implicitly) acts as a regularizer on the output embedding norm. In support of our argument that a subtoken attention distillation objective is superior to next-token prediction for learning new token embeddings, the combination via a sum of the two objectives (Token Distillation + NTP) generally yields worse results than Token Distillation alone. However, we can add a dynamic downweighting factor (see Section 4.2) to the next-token prediction objective (Token Distillation + NTP), which mostly alleviates the negative interference while keeping the regularizing effect.
@@
Vanilla NTP-tuning of all embeddings (denoted as tune all embeddings) yields disappointing results, even underperforming the subtoken mean initialization. We investigate this and observe a degradation of the original token embeddings. Note that for Token Distillation, we only optimize new token embeddings. Therefore, we also compare against the NTP baseline, where we similarly optimize only the new token embeddings. NTP outperforms NTP, supporting our analysis. However, even with this improvement, Token Distillation in turn still further surpasses NTP.
We note that masking gradient updates to the original embeddings during further training is not commonly done when adding new token embeddings, even during initial “embedding tuning” phases where all weights but the embedding layers are frozen. In these phases, the goal is to improve the initialized embeddings for new tokens or “adapt” the new embedding matrix to the existing Transformer layers (de Vries & Nissim, 2021). When normally sampling from a large-scale training corpus, the degradation of original embeddings will be less dramatic, as the number of new tokens per batch will be much lower. However, this is much less computationally efficient and – at least initially – still potentially harmful. We leave a further investigation of this as a practically useful direction for future work.
@@
ZeTT, which uses a pretrained hypernetwork, is the strongest baseline we compare against. We only compare against ZeTT when a pretrained hyper-network by Minixhofer et al. (2024) is available444Minixhofer et al. (2024) do not provide a checkpoint trained specifically for the -Instruct variant of Llama3-8B but show that a transfer from the base version is possible.. We first note that ZeTT yields quite impressive results, outperforming even our optimized NTP baseline. Evidently, the pretraining of ZeTT’s hypernetwork has internalized a better embedding prediction than the iterative gradient descent optimization on selected samples we employ for NTP. Nevertheless, our proposed method Token Distillation outperforms even ZeTT – without needing any hyper-network pretraining. It is important to note that at inference time, hyper-network based methods such as ZeTT are actually faster than Token Distillation (and NTP-based baselines), since only a single forward pass through the hyper-network per token is required. However, they require expensive pretraining for each new model and – in some cases – target domain. For example, ZeTT with a hypernetwork pretrained only on English naturally underperforms on language adaptation to other languages – as this is out of distribution from the hypernetwork training, whereas the multilingually pretrained variant available for one of the models performs more favorably (see Table 3). We note that in concurrent work, Minixhofer et al. (2025) have proposed to train embedding prediction hyper-networks with a distillation-based objective and show improvements over the next-token prediction objective, which is in line with our findings. However, this does not yet solve the domain-specificity problem of pretrained hyper-networks. Alternatively, we can actually use ZeTT-generated embeddings as a starting point and further tune them using our method for even better results. We analyze this possibility in Section 5.2.
@@ 5.2 In-Depth Analysis
Mistral-7B Llama3-8B Llama3-8B-i Llama3.1-8B Llama3.1-8B-i Llama3.2-3B Llama3.2-3B-i OLMo2-7B-i Avg. Original tokenization 93.4 88.1 99.2 91.1 97.8 65.2 90.9 94.0 90.0 Random (Hewitt, 2021) 0.0 0.0 0.0 0.0 0.0 0.0 0.2 0.0 0.0 Mean (Gee et al., 2022) 2.2 4.2 9.1 8.0 11.3 4.2 7.4 7.4 6.7 NTP (Lampinen et al., 2018) 34.0 54.3 70.6 49.7 71.6 29.0 53.9 70.0 54.1 ZeTT (Minixhofer et al., 2024) 15.3 25.4 28.4 – – – – – – Token Distillation 46.3 68.0 86.3 61.8 81.3 31.6 57.5 84.3 64.6 Token Distillation + NTP 48.3 68.0 85.3 60.2 79.7 38.2 63.4 80.9 65.5 Token Distillation + NTP 46.9 71.0 85.7 62.0 80.1 37.2 66.4 83.7 66.6
@@
In Table 4, we report LLM-as-a-Judge results on the correctness of generated definitions. This tests the limits of our method, as the new tokens are now more complicated, such as “Software Development” or even common phrases such as “spill the beans”. In this setting, even when using the original tokenization, the model sometimes is not able to generate correct definitions. In turn, the gap of Token Distillation to the original tokenization is also larger in this experiment. However, Token Distillation outperforms all other baselines by a very large margin. ZeTT (Minixhofer et al., 2024) in particular struggles in the multi-word token setup, likely because it is out-of-distribution from the hyper-network pretraining.
Mistral-7B Llama3-8B Llama3-8B-i Llama3.1-8B Llama3.1-8B-i Llama3.2-3B Llama3.2-3B-i OLMo2-7B-i Avg. Token Distillation (generated data) 62.4 67.1 67.6 67.0 70.1 57.1 62.3 60.9 64.3 Token Distillation (retrieved data) 62.8±0.5 67.3±0.2 67.6±0.3 67.3±0.5 71.0±0.2 56.2±1.9 63.1±0.2 61.2±0.3 64.6
@@
Choice of initial embedding for Token Distillation.
We find that the subtoken mean initialization as starting point yields better results than random initialization (see “Random + Token Distillation” in Table 7). If a pretrained embedding prediction hyper-network is available, we can instead use these predicted embeddings as the starting point. We run this experiment using ZeTT; this combination further improves performance compared to using the subtoken mean (see Table 7). We also compare refining the hypernetwork embeddings with NTP. This greatly improves over NTP but yields only minimal gains over using the hypernetwork embeddings themselves without further NTP.
Mistral-7B Llama3-8B Llama3-8B-i Llama3.1-8B Llama3.1-8B-i Llama3.2-3B Llama3.2-3B-i OLMo2-7B-i Avg. Token Distillation (Logits) 63.0±0.4 67.2±0.3 67.5±0.8 67.3±0.2 70.7±0.5 57.4±0.3 62.8±0.7 60.2±0.2 64.5 Token Distillation + NTP 63.0±0.5 67.2±0.3 66.7±1.6 67.2±0.3 70.6±0.4 57.6±0.3 62.2±0.3 59.8±0.5 64.3 Token Distillation + NTP 62.8±0.5 67.6±0.4 67.8±0.5 67.4±0.4 70.9±0.3 57.9±0.1 62.5±0.6 61.2±0.2 64.7 Token Distillation 62.8±0.5 67.3±0.2 67.6±0.3 67.3±0.5 71.0±0.2 56.2±1.9 63.1±0.2 61.2±0.3 64.6 Token Distillation (KL) + NTP 62.3±0.4 66.7±0.2 66.4±0.3 66.7±0.7 70.1±0.4 57.3±0.2 62.1±0.6 59.7±0.3 63.9 Token Distillation (KL) + NTP 62.8±0.5 67.1±0.4 66.3±1.2 67.1±0.2 70.5±0.4 57.5±0.5 62.5±0.2 60.5±0.5 64.3 Token Distillation (KL) 63.0±0.2 67.3±0.2 67.2±0.2 67.1±0.2 70.8±0.5 57.5±0.4 62.7±0.3 60.5±0.6 64.5
@@
Our main proposed method Token Distillation computes the distillation objective based on a mean-squared error (MSE) of last layer hidden states. A natural question is whether we can also use the “traditional” knowledge distillation objective (Hinton et al., 2015) based on the Kullback-Leibler divergence (KL-Div) (Kullback & Leibler, 1951). We analyze this and report the results in Table 6. We additionally include computing the MSE between logits instead of hidden states. For the KL-Div based results, we include different variations of combining the objective with a NTP-based objective. In general, the choice of distillation objective does not matter much. This is hardly surprising, as the logits are merely a linear projection (potentially with an additional normalization layer) of the hidden state; the log-probabilities used for KL-Div are then in turn merely an additional log-normalization via log-softmax. Combining the KL-based results with NTP-based objectives follows similar trends as their combinations with hidden state MSE-based objectives. We opt for the MSE on hidden states for two reasons: (1) In terms of implementation convenience, the objectives with a projection to logits over the vocabulary require a masking step for the “student” model, as the original “teacher” model does not have the new tokens in its vocabulary. (2) Choosing hidden states allows us to explore other layers than just the last one with potentially large computational speedups – we explore this next.
Mistral-7B Llama3-8B Llama3-8B-i Avg. Original Tokenization 64.5 69.8 70.6 68.3 Random (Hewitt, 2021) 56.7 58.7 60.6 58.6 Mean (Gee et al., 2022) 58.3 63.8 63.4 61.8 ZeTT (Minixhofer et al., 2024) 62.7 66.1 66.3 65.0 (Mean +) NTP 61.5 65.9 65.4 64.3 ZeTT + NTP 62.8 66.1 66.8 65.3 Random + Token Distillation 62.5 65.3 65.9 64.6 (Mean +) Token Distillation 63.5 67.3 68.0 66.3 ZeTT + Token Distillation 64.2 67.9 68.4 66.8 (Mean +) Token Distillation + NTP 63.2 67.2 68.4 66.3 ZeTT + Token Distillation + NTP 64.1 67.9 68.7 66.9
@@
In our main experiments, we apply the Token Distillation objective on the hidden states after the last layer. However, the last hidden state might not be the most optimal, as it might be overspecialized for next-token prediction (Rogers et al., 2020). In Figure 2, we plot the average performance on our biomedical benchmark datasets using Llama3-8B-Instruct while varying the target layer for extracting hidden states for our Token Distillation objective. Indeed we do see a slight downward trend as we approach the last layer – choosing a different layer than the last one can further boost our already strong results. We repeat the analysis with Llama3.1-8B-Instruct in Section A.2 with matching results.
Interestingly, good results can already be achieved with very early layers, e.g., after layer four. This suggests that much of the subtoken-specific contextual information is already added to the residual stream in early layers, which echoes findings of previous work (Vulić et al., 2020; Elhage et al., 2022; Lad et al., 2024; Nakash et al., 2025; Kaplan et al., 2025). This opens up an avenue to further speed up our method by terminating the forward pass after our target layer (and also saving that compute on the backward pass), which can be much faster if we select early target layers. We leave further exploration of this as an exciting direction for future work. For our main experiments, we choose to keep the last layer because this choice does not necessitate a model-specific sweep over target layers. Also, the last layer is a principled choice, as it guarantees that no subtoken interactions that are only modeled in later layers are excluded from the objective.